# Labelling Data with Unknown References

### v3

@@ Appendix E Rubrics
We present the full rubrics used in our experiments. Rubrics 2 and 3 have the specifications for the natural-language experiments. Both rubrics were passed verbatim to the LLM and used during annotation.
@@
Table 2: Rubric for the natural-language evaluation (IF) experiment. This rubric is meant to evaluate the quality of the output of an LLM. The rubric above was given to both the human annotators and the LLM from the No-Data Algorithm.
@@
Table 3: Rubric for the natural-language evaluation (OOF). This rubric was designed to be somewhat orthogonal to the IF version. That meant measuring a different problem altogether, albeit with some overlaps (e.g., part of c2b overlaps with c2a in IF, c2a here is c4 in IF). Some criteria are completely different, such as c1 and c5.
Appendix F Detailed Experimental Methods
For our first experiment, we used o3-mini (version 2025-01-31) with maximum tokens set to 50,000 and left all other parameters as default. GPT-4o (version 2024-05-13) and Qwen 2.5 VL 7B were called with maximum tokens set to 1,024 and temperature set to zero. DeepSeek (version r1-distill-qwen-32b) also had the temperature set to zero, but the maximum tokens set to 2,048 to account for the longer reasoning trace. For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters.
All LLMs were called via the Azure OpenAI API. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
F.1 Prompts
The prompts we used for our LLM evaluation on synthetic experiments are in Prompt 1 (evaluation) and Prompt 2 (generation). All exemplars are omitted for brevity. Whenever a parsing error occurred, we retried the call up to five times. If it failed on the fifth, we returned a random string. For the natural language experiments the prompts are in Prompt 4 (evaluator/verifier) and Prompt 5 (generator; part of the evaluator). These prompts required no exemplars. When parsing errors occurred, we retried up to five times, and otherwise returned a zero-labelled value.
You are labelling binary strings based on a rubric (given below).
First return the parts of the criterion that match the string, and the values.
Then return the label based on the aggregate function.
# Rubric:
<NATURAL-LANGUAGE RUBRIC GOES HERE>
# Aggregation Function:
Majority vote
Return your answer in the form:
|reasons|
(list of reasons)
|reasons|
|label|
(the label)
|label|
Table 1: System prompt for the evaluator (synthetic experiment). Anchors (‘|reasons|’, ‘|label|’) are used for parsing. The text contained within ‘|reasons|’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric.
You are given five datapoints and a set of criteria.
Based on these, pick the datapoint that matches the criteria exactly.
# Rubric:
<NATURAL-LANGUAGE RUBRIC GOES HERE>
# Datapoints:
<DATAPOINTS GO HERE>
Return your rationale, and then your chosen datapoint ONLY in the form:
|datapoint|
(the datapoint)
|datapoint|
@@ F.1 Prompts
You are a datapoint generator over binary strings.
Given a rubric (given below), a datapoint, and a label, return a similar datapoint that has the same
label, and fulfils the same conditions as the rubric.
For convenience, always start with the same datapoint: 0100. It will be easier to work with.
# Rubric:
<NL RUBRIC GOES HERE>
Return your rationale, and then the final datapoint in the form:
|datapoint|
(the datapoint)
|datapoint|
Table 3: System prompt for generation used in the ablation study from Appendix D.2. Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.
You are an LLM evaluator. You will be given a prompt and an response in West Frisian, meant for West Frisian readers.
Your job will be to verify if the response follows certain criteria and give a final binary score.
Check the output against the criteria below. If it fulfils the criteria, it should be a 1. Otherwise, 0.
If any of the criteria score a zero, the response must be zero.
# Criteria:
<CRITERIA GO HERE>
# Output format:
<OUTPUT FORMAT GOES HERE>
Table 4: System prompt for the natural-language evaluator and verifier. The output format requested is always in JSON. Depending on the setup, the output format and criteria will vary. For example, if it is meant to evaluate a single criterion, say, c2b, the definition will include only the definition of said criterion and the schema will specify solely ‘c2b’ as a key. Remark that the verifier will always expect the criteria from the IF rubric.
You are a paraphraser evaluating a prompt and an output for an LLM.
You will be given a datapoint (prompt/output), a label, and a list of reasons why that datapoint’s output has that label.
Your job will be to return a SIMILAR prompt and output, such that the OUTPUT (1) it matches the list of reasons, and (2) matches the label.
The output must match the values in the list of reasons.
Here’s the rubric used for these reasons:
<RUBRIC GOES HERE>
Your response must be in JSON using the following schema:
 "Prompt": the new, paraphrased user prompt.
 "Output": the new, paraphrased output fulfiling the criteria.
Only use the keys "Prompt" and "Output"
Table 5: System prompt for the natural-language generator. In this scenario the prompt will always include the rubric (either IF or OOF), as labelled by the evaluator itself.

### v4

@@ E.1 Rubrics
We present the full rubrics used in our experiments. Tables 10 and 11 have the specifications for the natural-language experiments. Both rubrics were used by the human annotators, and passed verbatim to the LLMs.
Table 10: Rubric for the natural-language evaluation (IP) experiment. This rubric is meant to evaluate the quality of the output of an LLM. The rubric above was given to both the human annotators and the LLM from the No-Data Algorithm.
@@
Table 11: Rubric for the natural-language evaluation setup (OOP). This rubric was designed to be mostly orthogonal to the IP version. That meant measuring a different problem altogether, albeit with some overlaps (e.g., part of c2b overlaps with c2a in IP, c2a here is c4 in IP). Some criteria are completely different, such as c1 and c5.
@@
E.2 Prompts
The prompts we used for our LLM evaluation on synthetic experiments are in Table 12 (evaluation) and Table 13 (generation). All exemplars are omitted for brevity. Whenever a parsing error occurred, we retried the call up to five times. If it failed on the fifth, we returned a random string. For the natural language experiments the prompts are in Table 15 (evaluator/verifier) and Table 16 (generator; part of the evaluator). These prompts required no exemplars. When parsing errors occurred, we retried up to five times, and otherwise returned a zero-labelled value.
Table 12: System prompt for the evaluator (synthetic experiment). Anchors (‘—reasons—’, ‘—label—’) are used for parsing. The text contained within ‘—reasons—’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric.
[…]
@@ E.2 Prompts
[…]
Table 14: System prompt for generation used in the ablation study from Appendix D.3. Unlike Prompt LABEL:pro:generationprompt, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.
[…]
Table 15: System prompt for the natural-language evaluator and verifier. The output format requested is always in JSON. Depending on the setup, the output format and criteria will vary. For example, if it is meant to evaluate a single criterion, say, c2b, the definition will include only the definition of said criterion and the schema will specify solely ‘c2b’ as a key. Remark that the verifier will always expect the criteria from the IP rubric.
[…]
Table 16: System prompt for the natural-language generator. In this scenario the prompt will always include the rubric (either IP or OOP), as labelled by the evaluator itself.
[…]